- Determine how many solutions exist
- Use either elimination or substitution to find the solutions (if any)
- Graph the two lines, labeling the x-intercepts, y-intercepts, and points
of intersection
y = 2x + 3 and y = -x – 4

Answer by checkley77(12844)

y = 2x + 3
y = -x – 4 NOW SUBTRACT.
----------------------------
0=3X+7
3X=-7
X=-7/3 ANSWER.
Y=-(-7/3)-4
Y=7/3-4
Y=(7-12)/3
Y=-5/3 ANSWER.
PROOF:
-5/3=2*-7/3+3
-5/3=-14/3+3
-5/3=(-14+9)/3
-5/3=-5/3